# WOD-E2E: Waymo Open Dataset for End-to-End Driving in Challenging Long-tail Scenarios

### v2

@@ 1 Introduction
Autonomous driving systems have traditionally followed a modular design approach that decomposes the driving task into distinct sub-tasks such as perception, prediction, and planning  [38, 13, 15, 27, 36, 23]. While this modular design offers benefits in terms of interpretability and debugging, the research community has recently shifted its attention to exploring vision-based end-to-end (E2E) architectures [34, 22, 30, 7, 37]. This shift is primarily driven by the inherent scalability of E2E systems, which directly map raw sensor data to driving actions, reducing the underlying system complexity and the need for rater annotations of intermediate concepts [33]. Furthermore, as previous works [14, 28] indicate, there is a promise of leveraging multi-modal large language models (MLLMs) and their world knowledge for E2E driving.
@@
Despite this promise, current real-world E2E driving datasets, such as NAVSIM [8], WOMD [10] and CoVLA [1], predominantly feature nominal driving scenarios that do not fully expose systems to the long tail of possible real-world situations. This scarcity of long-tail examples hinders the accurate evaluation of the true potential, robustness, and generalization ability of E2E driving systems.
@@
In this paper, we introduce the newly released Waymo Open Dataset for End-to-End Driving (WOD-E2E), which explicitly focuses on long-tail situations. As shown in Figure 1, WOD-E2E features rare real-world scenarios, which occur with a frequency of less than 0.03%. We provide 4,021 challenging driving segments comprising approximately 12 hours in total, where each segment contains 8 surrounding cameras covering a 360-degree field of view, high-level routing information, ego vehicle position history, and 5s of its future trajectory. These driving segments are collected from a mixture of autonomous and manual driving. Moreover, we observe that previous open-loop metrics often fail to adequately evaluate the driving performance in these long-tail scenarios. The popular Average Distance Error (ADE) or L2 error metric captures only the error between a prediction and a single future ground truth trajectory, despite the driving behavior being inherently multi-modal, where multiple reasonable future trajectories are possible. Predictive metrics, such as PDMS scores [8], require annotated positions and future trajectories of road agents to calculate collision rates, and thus become impractical in many long-tail scenarios involving novel or hard-to-detect objects (e.g., the flock of birds shown in Figure 1). Furthermore, off-road behaviors typically incur high penalties in PDMS, yet in numerous safety-critical long-tail scenarios, an autonomous vehicle might reasonably deviate partially off-road to avoid an emergency. To address these limitations, WOD-E2E dataset also includes a subset of human driving preference labels, providing expert ratings on multiple potential trajectories in each example. Leveraging these labels, we propose a novel open-loop evaluation metric, the Rater Feedback Score (RFS), to better evaluate the E2E driving performance in an open-loop setting.
@@
We conduct rigorous studies with robust baseline models to verify the dataset and RFS. Since the dataset release, we have garnered significant interest from the research community, with numerous methods already submitted and evaluated on our public leaderboard. The diversity of these top-performing methods, employing approaches such as MLLMs [21, 29], diffusion models [17], and CNN/ViT with GRU/MLP architectures [24], further underscores the utility of the WOD-E2E dataset and its promise to drive further advances in end-to-end autonomous driving research. Our contribution can be summarized as:
@@ 2.1 End-to-end autonomous driving research
The paradigm of E2E autonomous driving, directly mapping raw sensor inputs to control outputs, continues to be a vibrant area of research, seeking to overcome the complexities in traditional modular pipelines [14, 33]. Recent works have significantly advanced the capabilities of E2E systems, particularly through the use of foundation models. Overall, the current methods can be divided into three categories:
Bird’s-Eye-View (BEV) Based E2E Planner: This type of method aims to fuse information from multiple sensors into a single, comprehensive BEV representation, from which both perception and planning tasks can be directly performed. UniAD [12] exemplifies this by propagating BEV queries from its perception module to downstream tasks such as tracking, motion forecasting, and occupancy prediction, ultimately enabling end-to-end planning. Similarly, BEV-Planner [16] focuses on learning an explicit planning policy directly from BEV features, demonstrating how dense BEV representations can facilitate robust end-to-end control. These approaches move beyond explicit intermediate perception outputs for planning. Overall, these unified BEV-centric methods offer advantages in terms of computational efficiency and coherence by providing a consistent spatial understanding across various driving sub-tasks.
Multi-modal Large Language Model Based E2E Planner: A prominent trend involves leveraging Multimodal Large Language Models (MLLMs) to imbue E2E driving systems with enhanced reasoning capabilities and world knowledge. DriveGPT4 [37] utilizes LLMs to both explain vehicle actions and predict control signals in an iterative question-and-answer format. DriveVLM [28] applies chain-of-thought for end-to-end driving, while VLP [22] applies the reasoning of MLLMs directly on the Bird’s-Eye-View (BEV) space. EMMA [14] leverages Gemini to process multiple driving tasks, including planning, 3D detection, and road understanding, within a unified language space. OpenEMMA [34] and LightEMMA [25] follow a similar paradigm to build an open-source and lightweight version, respectively. Additionally, S4-Driver [33] proposes to lift the vision tokens from MLLMs to a 3D space.
Diffusion Based E2E Planner: Diffusion models excel at capturing the multi-modal nature of driving actions and generating diverse, plausible trajectories. Notably, DiffusionDrive [17] introduces a truncated diffusion policy and efficient cascade decoder for real-time E2E driving. EnDfuser [32] further explores using diffusion ensembles to estimate uncertainty in trajectory planning, leveraging fused camera and LiDAR features to produce distributions of candidate trajectories.
@@ 2.2 End-to-end autonomous driving open dataset
A multitude of autonomous driving datasets are available today, supporting a diverse range of driving tasks. Notable examples include Kitti [11], Argoverse [5], Argoverse 2 [31], WOD-Perception [26], and V2V4Real [35]. While these datasets serve various purposes, a specific subset focuses on end-to-end driving. Among the most prominent open datasets in this category are nuScenes [2], NAVSIM [8], WOMD [10], and CoVLA [1].
@@
nuScenes [2] is initially developed for perception tasks and features multiple sensor modalities. While recent research [28, 14, 22] has explored end-to-end driving on this dataset, often using ADE as a primary performance indicator, the core focus of nuScenes remains perception rather than planning. Some studies [16, 39] have observed that even simple extrapolation of historical behavior can yield strong performance without relying on camera images, suggesting that nuScenes may not be ideally suited for complex planning tasks.
@@ 2.2.2 NAVSIM
NAVSIM [8] is a compact simulation and benchmarking framework built upon a filtered version of nuPlan [3]. Its core contribution lies in enabling large-scale real-world evaluation through a non-reactive simulator, which effectively bridges the gap between open-loop and closed-loop testing via simulation-based metrics. While NAVSIM has helped to significantly advance end-to-end driving research, its approach presents two major limitations that motivated our work. First, as a simulation framework, it relies on filtering existing datasets rather than providing a raw data collection effort specifically for long-tail events, which may preclude it from capturing the full, nuanced diversity of real-world long-tail scenarios. Second, the PDMS proposed in NAVSIM—which heavily prioritizes ego progress and comfort, along with time-to-collision (TTC)—may prove insufficient for true safety-critical situations. For instance, TTC is challenging to measure with amorphous obstacles like a flock of birds, as depicted in Figure 1, and the metric of comfort should be secondary to safety when the vehicle must perform an emergency maneuver, such as avoiding a falling scooter (Figure 1). Our dataset and evaluation methodology are explicitly designed to overcome these two limitations by providing targeted, diverse, long-tail data and a more safety-focused scoring mechanism.
@@ 2.2.3 WOMD
The Waymo Open Motion Dataset (WOMD) [10] is a component of the broader Waymo Open Dataset, with a specific emphasis on motion prediction and behavior research. Although recent works such as MoST [19] and S4-Driver [33] conduct end-to-end driving research on it, WOMD is primarily designed for motion prediction and for modeling complex agent interactions, rather than planning. Furthermore, the lack of full camera images (only embeddings are provided) makes it difficult for external researchers to conduct comprehensive E2E research.
@@ 2.2.4 CoVLA
CoVLA [1] provides a large-scale, richly annotated collection of real-world driving scenarios, integrating vision, language, and action modalities. It is designed to enable the training of Vision-Language-Action models that can generate descriptive scene captions and predict vehicle trajectories. While CoVLA’s automated captioning aims for diversity and covers a wide range of common driving conditions, the available information does not detail specific mechanisms for over-sampling or synthesizing rare, safety-critical long-tail events beyond general diversity.
@@ 3.1.2 Camera Data
This dataset includes images from eight cameras, providing 360-degree coverage around the vehicle: front, front left, front right, side left, side right, rear, rear left, and rear right. The sensor layout configuration is similar to that described in [26]. For each direction, a single JPEG image is provided. Alongside the image data, we supply camera intrinsics and extrinsics, which define the camera’s internal parameters and its position relative to the vehicle’s center, respectively. These parameters enable the projection of 3D trajectories onto the camera images. Each driving segment includes 10Hz camera video sequences. Training data spans 20 seconds, while testing data covers 12 seconds, with the subsequent 8 seconds of future data hidden for evaluation purposes.
@@ 3.1.3 Routing Information
We provide a routing input for the model in the form of a high-level command, following conventional academic benchmarks [12, 4].
@@ 3.2 Quantitative Rareness Comparison
In this section, we quantitatively compare the rarity of WOD-E2E against other popular E2E driving datasets. To achieve a standardized, impartial rarity assessment, we utilized a large language model, Gemini 2.5 Pro [6], to score the test set of each dataset. The model was provided with the front camera sequences and a detailed scoring prompt outlining four tiers of rarity based on complexity, risk, and long-tail factors. The prompt required the output to be a JSON object containing the rarity_score that is ranged from 0-100, identified rare_factors, and a reasoning trace for maximum transparency.
@@ 3.4.2 Trajectory Sampling
Trajectory sampling is the process of generating a diverse set of possible motion plans for later human review and selection in a specific driving scenario. Our approach utilizes an existing machine learning model, such as Wayformer [20], to produce an initial set of up to 64 diverse trajectories for a given critical moment. These trajectories are generated using various inputs, including perception detections, mapping elements, and predicted behaviors of other road agents.
@@ 3.5 Rater Feedback Score
The base thresholds follows WOMD [10], and they are set as $\bar{\tau}_{\mathrm{lat}}=1.0,\bar{\tau}_{\mathrm{lng}}=4.0$ at $t=3$ and $\bar{\tau}_{\mathrm{lat}}=1.8,\bar{\tau}_{\mathrm{lng}}=7.2$ at $t=5$, where the longitudinal threshold $\bar{\tau}_{\mathrm{lng}}$ is always set to be 4 times larger than the lateral threshold $\bar{\tau}_{\mathrm{lat}}$. These base thresholds are scaled based on the initial speed $v$ (m/s) of the rater trajectory. The scaling function is a piece-wise linear function of $v$:
@@
Intuitively, we assign either the flat score $s_{\mathrm{rater}}$, if a predicted trajectory is within the trust region, or the score exponentially decayed from $s_{\mathrm{rater}}$. Then, the final score is determined by choosing the maximum score over all rater specified trajectories, followed by averaging over $t=3,5$ and flooring with 44.
@@ 4.1 Baseline Model Setup
We use a highly simplified version of EMMA [14], which we call NaiveEMMA, as our baseline model. The architecture of NaiveEMMA is illustrated in Figure 9. NaiveEMMA is finetuned directly from Gemini Flash [6] and has not been trained on any internal driving datasets: it is finetuned exclusively on the released WOD-E2E training split. The model consumes a combined image from all eight cameras at the current timestep, concatenated into a single $768\times 768$ resolution image. It also takes in 3 seconds of past ego-status history and the high-level routing input. Crucially, it does not use past camera frames. Note that NaiveEMMA omits several advanced components of the original EMMA model, specifically generalist task training mixtures, Chain-of-Thought reasoning, and any test-time scaling methods.
Figure 9: Architecture of NaiveEMMA, which serves as the challenge leaderboard baseline. NaiveEMMA is a highly simplified version of EMMA [14], fine-tuned from Gemini Flash [6]. The model takes as input all 8 camera images, 3 seconds of past ego-status history, and the high-level routing input. It then predicts the future trajectory in 5 seconds

### v3

@@ 1 Introduction
Autonomous driving systems have traditionally followed a modular design approach that decomposes the driving task into distinct sub-tasks such as perception, prediction, and planning  [yurtsever2020survey, hwang2022cramnet, li2022bevformer, sun2022swformer, xu2025v2xvit2, pan2024clip]. While this modular design offers benefits in terms of interpretability and debugging, the research community has recently shifted its attention to exploring vision-based end-to-end (E2E) architectures [xing2025openemma, pan2024vlp, wang2025adawm, cui2025vilad, xu2024drivegpt4]. This shift is primarily driven by the inherent scalability of E2E systems, which directly map raw sensor data to driving actions, reducing the underlying system complexity and the need for rater annotations of intermediate concepts [xie2025s4]. Furthermore, as previous works [hwang2025emma, tian2024drivevlm] indicate, there is a promise of leveraging multi-modal large language models (MLLMs) and their world knowledge for E2E driving.
@@
Despite this promise, current real-world E2E driving datasets, such as NAVSIM [Dauner2024NEURIPS], WOMD [ettinger2021large] and CoVLA [arai2025covla], predominantly feature nominal driving scenarios that do not fully expose systems to the long tail of possible real-world situations. This scarcity of long-tail examples hinders the accurate evaluation of the true potential, robustness, and generalization ability of E2E driving systems.
@@
In this paper, we introduce the newly released Waymo Open Dataset for End-to-End Driving (WOD-E2E), which explicitly focuses on long-tail situations. As shown in Figure 1, WOD-E2E features rare real-world scenarios, which occur with a frequency of less than 0.03%. We provide 4,021 challenging driving segments comprising approximately 12 hours in total, where each segment contains 8 surrounding cameras covering a 360-degree field of view, high-level routing information, ego vehicle position history, and 5s of its future trajectory. These driving segments are collected from a mixture of autonomous and manual driving. Moreover, we observe that previous open-loop metrics often fail to adequately evaluate the driving performance in these long-tail scenarios. The popular Average Distance Error (ADE) or L2 error metric captures only the error between a prediction and a single future ground truth trajectory, despite the driving behavior being inherently multi-modal, where multiple reasonable future trajectories are possible. Predictive metrics, such as PDMS scores [Dauner2024NEURIPS], require annotated positions and future trajectories of road agents to calculate collision rates, and thus become impractical in many long-tail scenarios involving novel or hard-to-detect objects (e.g., the flock of birds shown in Figure 1). Furthermore, off-road behaviors typically incur high penalties in PDMS, yet in numerous safety-critical long-tail scenarios, an autonomous vehicle might reasonably deviate partially off-road to avoid an emergency. To address these limitations, WOD-E2E dataset also includes a subset of human driving preference labels, providing expert ratings on multiple potential trajectories in each example. Leveraging these labels, we propose a novel open-loop evaluation metric, the Rater Feedback Score (RFS), to better evaluate the E2E driving performance in an open-loop setting.
@@
We conduct rigorous studies with robust baseline models to verify the dataset and RFS. Since the dataset release, we have garnered significant interest from the research community, with numerous methods already submitted and evaluated on our public leaderboard. The diversity of these top-performing methods, employing approaches such as MLLMs [pal2025poutinevisionlanguagetrajectorypretraining, wang2025hmvlm], diffusion models [liao2025diffusiondrive], and CNN/ViT with GRU/MLP architectures [ParkSwinTrajectoryTR], further underscores the utility of the WOD-E2E dataset and its promise to drive further advances in end-to-end autonomous driving research. Our contribution can be summarized as:
@@ 2.1 End-to-end autonomous driving research
The paradigm of E2E autonomous driving, directly mapping raw sensor inputs to control outputs, continues to be a vibrant area of research, seeking to overcome the complexities in traditional modular pipelines [hwang2025emma, xie2025s4]. Recent works have significantly advanced the capabilities of E2E systems, particularly through the use of foundation models. Overall, the current methods can be divided into three categories:
Bird’s-Eye-View (BEV) Based E2E Planner: This type of method aims to fuse information from multiple sensors into a single, comprehensive BEV representation, from which both perception and planning tasks can be directly performed. UniAD [hu2023uniad] exemplifies this by propagating BEV queries from its perception module to downstream tasks such as tracking, motion forecasting, and occupancy prediction, ultimately enabling end-to-end planning. Similarly, BEV-Planner [li2024bevplanner] focuses on learning an explicit planning policy directly from BEV features, demonstrating how dense BEV representations can facilitate robust end-to-end control. These approaches move beyond explicit intermediate perception outputs for planning. Overall, these unified BEV-centric methods offer advantages in terms of computational efficiency and coherence by providing a consistent spatial understanding across various driving sub-tasks.
Multi-modal Large Language Model Based E2E Planner: A prominent trend involves leveraging Multimodal Large Language Models (MLLMs) to imbue E2E driving systems with enhanced reasoning capabilities and world knowledge. DriveGPT4 [xu2024drivegpt4] utilizes LLMs to both explain vehicle actions and predict control signals in an iterative question-and-answer format. DriveVLM [tian2024drivevlm] applies chain-of-thought for end-to-end driving, while VLP [pan2024vlp] applies the reasoning of MLLMs directly on the Bird’s-Eye-View (BEV) space. EMMA [hwang2025emma] leverages Gemini to process multiple driving tasks, including planning, 3D detection, and road understanding, within a unified language space. OpenEMMA [xing2025openemma] and LightEMMA [qiao2025lightemma] follow a similar paradigm to build an open-source and lightweight version, respectively. Additionally, S4-Driver [xie2025s4] proposes to lift the vision tokens from MLLMs to a 3D space.
Diffusion Based E2E Planner: Diffusion models excel at capturing the multi-modal nature of driving actions and generating diverse, plausible trajectories. Notably, DiffusionDrive [liao2025diffusiondrive] introduces a truncated diffusion policy and efficient cascade decoder for real-time E2E driving. EnDfuser [wintel2025using] further explores using diffusion ensembles to estimate uncertainty in trajectory planning, leveraging fused camera and LiDAR features to produce distributions of candidate trajectories.
@@ 2.2 End-to-end autonomous driving open dataset
A multitude of autonomous driving datasets are available today, supporting a diverse range of driving tasks. Notable examples include Kitti [geiger2013vision], Argoverse [chang2019argoverse], Argoverse 2 [wilson2023argoverse], WOD-Perception [sun2020scalability], and V2V4Real [xu2023v2v4real]. While these datasets serve various purposes, a specific subset focuses on end-to-end driving. Among the most prominent open datasets in this category are nuScenes [caesar2020nuscenes], NAVSIM [Dauner2024NEURIPS], WOMD [ettinger2021large], and CoVLA [arai2025covla].
@@
nuScenes [caesar2020nuscenes] is initially developed for perception tasks and features multiple sensor modalities. While recent research [tian2024drivevlm, hwang2025emma, pan2024vlp] has explored end-to-end driving on this dataset, often using ADE as a primary performance indicator, the core focus of nuScenes remains perception rather than planning. Some studies [li2024bevplanner, zhai2023rethinking] have observed that even simple extrapolation of historical behavior can yield strong performance without relying on camera images, suggesting that nuScenes may not be ideally suited for complex planning tasks.
@@ 2.2.2 NAVSIM
NAVSIM [Dauner2024NEURIPS] is a compact simulation and benchmarking framework built upon a filtered version of nuPlan [caesar2021nuplan]. Its core contribution lies in enabling large-scale real-world evaluation through a non-reactive simulator, which effectively bridges the gap between open-loop and closed-loop testing via simulation-based metrics. While NAVSIM has helped to significantly advance end-to-end driving research, its approach presents two major limitations that motivated our work. First, as a simulation framework, it relies on filtering existing datasets rather than providing a raw data collection effort specifically for long-tail events, which may preclude it from capturing the full, nuanced diversity of real-world long-tail scenarios. Second, the PDMS proposed in NAVSIM—which heavily prioritizes ego progress and comfort, along with time-to-collision (TTC)—may prove insufficient for true safety-critical situations. For instance, TTC is challenging to measure with amorphous obstacles like a flock of birds, as depicted in Figure 1, and the metric of comfort should be secondary to safety when the vehicle must perform an emergency maneuver, such as avoiding a falling scooter (Figure 1). Our dataset and evaluation methodology are explicitly designed to overcome these two limitations by providing targeted, diverse, long-tail data and a more safety-focused scoring mechanism.
@@ 2.2.3 WOMD
The Waymo Open Motion Dataset (WOMD) [ettinger2021large] is a component of the broader Waymo Open Dataset, with a specific emphasis on motion prediction and behavior research. Although recent works such as MoST [mu2024most] and S4-Driver [xie2025s4] conduct end-to-end driving research on it, WOMD is primarily designed for motion prediction and for modeling complex agent interactions, rather than planning. Furthermore, the lack of full camera images (only embeddings are provided) makes it difficult for external researchers to conduct comprehensive E2E research.
@@ 2.2.4 CoVLA
CoVLA [arai2025covla] provides a large-scale, richly annotated collection of real-world driving scenarios, integrating vision, language, and action modalities. It is designed to enable the training of Vision-Language-Action models that can generate descriptive scene captions and predict vehicle trajectories. While CoVLA’s automated captioning aims for diversity and covers a wide range of common driving conditions, the available information does not detail specific mechanisms for over-sampling or synthesizing rare, safety-critical long-tail events beyond general diversity.
@@ 3.1.2 Camera Data
This dataset includes images from eight cameras, providing 360-degree coverage around the vehicle: front, front left, front right, side left, side right, rear, rear left, and rear right. The sensor layout configuration is similar to that described in [sun2020scalability]. For each direction, a single JPEG image is provided. Alongside the image data, we supply camera intrinsics and extrinsics, which define the camera’s internal parameters and its position relative to the vehicle’s center, respectively. These parameters enable the projection of 3D trajectories onto the camera images. Each driving segment includes 10Hz camera video sequences. Training data spans 20 seconds, while testing data covers 12 seconds, with the subsequent 8 seconds of future data hidden for evaluation purposes.
@@ 3.1.3 Routing Information
We provide a routing input for the model in the form of a high-level command, following conventional academic benchmarks [hu2023uniad, casas2021mp3].
@@ 3.2 Quantitative Rareness Comparison
In this section, we quantitatively compare the rarity of WOD-E2E against other popular E2E driving datasets. To achieve a standardized, impartial rarity assessment, we utilized a large language model, Gemini 2.5 Pro [comanici2025gemini], to score the test set of each dataset. The model was provided with the front camera sequences and a detailed scoring prompt outlining four tiers of rarity based on complexity, risk, and long-tail factors. The prompt required the output to be a JSON object containing the rarity_score that is ranged from 0-100, identified rare_factors, and a reasoning trace for maximum transparency.
@@ 3.4.2 Trajectory Sampling
Trajectory sampling is the process of generating a diverse set of possible motion plans for later human review and selection in a specific driving scenario. Our approach utilizes an existing machine learning model, such as Wayformer [nayakanti2023wayformer], to produce an initial set of up to 64 diverse trajectories for a given critical moment. These trajectories are generated using various inputs, including perception detections, mapping elements, and predicted behaviors of other road agents.
@@ 3.5 Rater Feedback Score
The base thresholds follows WOMD [ettinger2021large], and they are set as $\bar{\tau}_{\mathrm{lat}}=1.0,\bar{\tau}_{\mathrm{lng}}=4.0$ at $t=3$ and $\bar{\tau}_{\mathrm{lat}}=1.8,\bar{\tau}_{\mathrm{lng}}=7.2$ at $t=5$, where the longitudinal threshold $\bar{\tau}_{\mathrm{lng}}$ is always set to be 4 times larger than the lateral threshold $\bar{\tau}_{\mathrm{lat}}$. These base thresholds are scaled based on the initial speed $v$ (m/s) of the rater trajectory. The scaling function is a piece-wise linear function of $v$:
@@
Intuitively, we assign either the flat score $s_{\mathrm{rater}}$, if a predicted trajectory is within the trust region, or the score exponentially decayed from $s_{\mathrm{rater}}$. Then, the final score is determined by choosing the maximum score over all rater specified trajectories, followed by averaging over $t=3,5$ and flooring with $4$.
@@ 4.1 Baseline Model Setup
We use a highly simplified version of EMMA [hwang2025emma], which we call NaiveEMMA, as our baseline model. The architecture of NaiveEMMA is illustrated in Figure 9. NaiveEMMA is finetuned directly from Gemini Flash [comanici2025gemini] and has not been trained on any internal driving datasets: it is finetuned exclusively on the released WOD-E2E training split. The model consumes a combined image from all eight cameras at the current timestep, concatenated into a single $768\times 768$ resolution image. It also takes in 3 seconds of past ego-status history and the high-level routing input. Crucially, it does not use past camera frames. Note that NaiveEMMA omits several advanced components of the original EMMA model, specifically generalist task training mixtures, Chain-of-Thought reasoning, and any test-time scaling methods.
Figure 9: Architecture of NaiveEMMA, which serves as the challenge leaderboard baseline. NaiveEMMA is a highly simplified version of EMMA [hwang2025emma], fine-tuned from Gemini Flash [comanici2025gemini]. The model takes as input all 8 camera images, 3 seconds of past ego-status history, and the high-level routing input. It then predicts the future trajectory in 5 seconds